You have 7 identical Probability books, 5 identical chocolate bars, and 2 identical Minivans to distribute

to your friends Eric, Anna, and Katie. In how many ways can you distribute the gifts among your

friends?
Answer #1

let x1,y1,z1 are probability books,chocolate bars and minivans goes to Eric

x2,y2,z2 are probability books,chocolate bars and minivans goes to Anna

x3,y3,z3 are probability books,chocolate bars and minivans goes to Katie

therefore x1+x2+x3 =7

for which number of ways =(7+3-1C3-1) =(9C2) =36

y1+y2+y3 =5

for which number of ways =(5+3-1C3-1) =(7C2) =21

and z1+z2+z3 =2

for which number of ways =(2+3-1C3-1) =(4C2) =6

total number of ways =36*21*6 =4536